OpenStudy (avai70178):

Can someone show me how to find the zeroes of this function? (A.K.A. x intercepts of the function) f(x) = x^3 + 2x^2 - 4x - 8

ganeshie8 (ganeshie8):

Have you tried factoring ?

ganeshie8 (ganeshie8):

f(x) = x^3 + 2x^2 - 4x - 8 group first two terms group last two terms

OpenStudy (avai70178):

okay so (x^3+2x^2) -(4x-8)?

OpenStudy (avai70178):

oh okay i see what happened there

ganeshie8 (ganeshie8):

careful when you put a parenthesis after a minus sign what you're essentially doing is factoring out a "-1"

ganeshie8 (ganeshie8):

f(x) = x^3 + 2x^2 - 4x - 8 = (x^3+2x^2) -1(4x + 8)

ganeshie8 (ganeshie8):

Next, factor the GCF from each of the groups

OpenStudy (avai70178):

x^2(x+2)

OpenStudy (avai70178):

how would i do the other half?

ganeshie8 (ganeshie8):

you may factor out 4

ganeshie8 (ganeshie8):

f(x) = x^3 + 2x^2 - 4x - 8 = (x^3+2x^2) -1(4x + 8) = x^2(x+2) - 4(x+2)

ganeshie8 (ganeshie8):

Notice that both have "x+2" in common factor it out

OpenStudy (marcelie):

you can do this too just enter that whole equation to your calculator.

OpenStudy (anonymous):

The roots (zeros) are the x values where the graph intersects the x-axis. To find the roots (zeros), replace y with 0 and solve for x. x=−2,2